# Scalable Bayesian Inference for Bradley–Terry Models with Ties: An Application to Honour Based Abuse

### v1

@@ Abstract
Honour based abuse covers a wide range of family abuse including female genital mutilation and forced marriage. Safeguarding professionals need to identify where abuses are happening in their local community to best support those at risk of these crimes and take preventative action. However, there is little local data about these kinds of crime. To tackle this problem, we ran comparative judgement surveys to map abuses at local level. In previous comparative judgement studies, participants reported fatigue associated with comparisons between areas with similar levels of abuse. Allowing for ties reduces fatigue, but increase the computational complexity when fitting the model. We designed an efficient Markov Chain Monte Carlo algorithm to fit the model, allowing for a wide range of prior distributions on the model parameters. Working with South Yorkshire Police and Oxford Against Cutting, we mapped the risk of honour based abuse at community level in two counties in the UK.
@@ 1 Introduction
Comparative judgement models have been to used to analyse a wide range of problems from predicting the outcomes of football leagues Cattelan et al., (2012) to student assessment Pollitt, (2012). The aim of a comparative judgement study is to estimate the qualities of a set of objects. By allowing study participants to compare objects in terms of the quantity of interest rather than give absolute assessments of that quantity, comparative judgement models can elicit reliable and informative responses. The Bradley–Terry (BT) model Bradley and Terry, (1952) is perhaps the most common model for analysing comparative judgement data. It used to rank and estimate features of a set of objects based on pair-wise comparisons. It assigns a quality parameter to each object in the study and estimates the value of these parameters from pairwise comparisons.
@@
We were motivated by a project to tackle Honour Based Abuse in the UK and support the development of community level safeguarding strategies. The specific aim of our project was to map the prevalence of honour based abuse at local level in two counties in the UK. Abuse to control family members in order to protect the family’s honour or standing is called honour based abuse. Perpetrators of honour based abuse often perceive a victim to be putting the family’s honour at risk, or to be bringing shame upon the family Aplin, (2019). Examples include Female Genital Mutilation, where a women or girl’s genitals are mutilated in order to make her more honourable, or forced marriage, where individuals are coerced to marry someone who is perceived to be honourable. These abuse causes significant physical and emotional harm to the victim and can cause both immediate and long-term health issues.
@@
To map the prevalence of abuse at community level, we considered analysing existing data. However, we found existing data to be insufficient. Currently, the UK Government only provides regional level data about forced marriage, and national level data about other types of honour based abuse Forced Marriage Unit, (2019). This makes developing effective safeguarding strategies for specific communities difficult. Other local level data does exist, as those providing services to victims may collect their own data. However, victims may be supported by a number of independent services, for example, law enforcement, health professionals or women’s refuges (Aplin,, 2019, §$\S 7$). This means that no one service has the complete picture of where honour based abuse is happening. Accessing data from independent services would require a large number of data sharing agreements and the transfer of sensitive data about victims of abuse, which was not feasible. Thus, we decided to collect our data using a method that avoids collecting personally identifying information.
@@
We decided to carry out a comparative judgement study. A comparative judgement study works by showing study participants pairs of wards in the community and asking them to choose which ward has the higher rate of the abuse in question. Study participants, known as judges, should be people who are familiar with the community and abuse, however the wisdom of the crowd type approach means that judges do not need to have a perfect knowledge. Running a comparative judgement study is a viable and effective way to measure the rate of abuses at community level, without the need to collect personally identifying or sensitive data. Indeed, it has been previously used to map forced marriage at community level in Nottinghamshire Seymour et al., 2022a and deprivation at sub-ward level in Dar es Salaam, Tanzania Seymour et al., 2022b . However, in previous studies, judges reported fatigue and difficulty providing pairwise comparisons when the levels of abuse in pairs of wards were similar Seymour et al., 2022a . In particular judges, found it difficult to repeatedly make comparisons about pairs of wards which both had low levels of abuse. This has also been found to be an issue when using adaptive comparative judgement study designs Jones and Davies, (2024). To mitigate this issue in our study, we allowed the judges to say (i) the level of abuse in one ward was higher than the other, (ii) the level of abuse was similar – or tied – in both wards, or (iii) they wanted to skip that particular pair of wards.
@@
BT models allowing for tied comparisons are well developed, with the model described in Rao and Kupper, (1967) commonly used. The outcome of each comparison is modelled by a multinomial distribution, with the events being a win, a loss, and a tie. Inference for models with ties is, however, limited. Inference methods for simple models allowing for tied comparisons are fast and efficient but require large amounts of data. Inference methods have been developed for more sophisticated models can reduce the amount of data that need to be collected, but are slow to and expensive to use. As our application is one with a limited number of judges, the statistical aim of our project was to develop a method that allowed for correlated prior distributions, but given the large number of model parameters, we also required a scalable method. We developed a new representation of the tied model by introducing a series of Pólya-Gamma latent variables. This allows for both sophisticated prior correlation between the model parameters and scalable inference. or computing maximum likelihood estimates, an iterative algorithm is developed in Rao and Kupper, (1967). Two Markov Chain Monte Carlo methods exist for generating samples from posterior distributions relevant to this model. In Caron and Doucet, (2012), the authors introduce a series of exponentially distribution latent variables to perform scalable inference. However, this method requires independent prior distributions to be placed on the model parameters, and an improper prior distribution on the parameter destining the importance of ties. In contrast, in Seymour et al., 2022b , the authors propose a Metropolis-Hastings Random Walk algorithm that allows for a wide range of correlations between the model parameters, and a choice of any prior distribution on the tie parameter, but scales poorly with respect to the number of wards in the study.
@@ 1.1 Data
In the county of Oxfordshire, we worked with Oxford Against Cutting, a charity tackling FGM and honour based abuse in the county. We focused on the city of Oxford and town of Banbury in this study, and judges were shown pairs of wards in these two locations and asked which had a higher rate of honour based abuse. Judges could also say the wards had a similar level of abuse.
@@ 2.1 The Standard Bradley–Terry Model
One of the most widely used models for modelling comparative judgement data is the-so-called BT model Bradley and Terry, (1952) and is defined as follows. Consider a set of $N$ wards. We assign a parameter $\lambda_{i}\in\mathbb{R}$ to each ward describing its quality. Large positive values of $\lambda_{i}$ correspond to wards with low prevelance of honour based abuse and large negative values to wards with high prevalence of abuse. When comparing wards $i$ and $j$, the probability ward $i$ is judged to be superior to ward $j$ depends on the difference in quality parameters
$\textrm{logit}(\pi_{ij})=\lambda_{i}-\lambda_{j}\iff\pi_{ij}=\frac{\exp(\lambda_{i})}{\exp(\lambda_{i})+\exp(\lambda_{j})}\qquad(i\neq j,1\leq i,j\leq N).$ (1)
Let $n_{ij}$ be the number of times wards $i$ and $j$ are compared, $y_{ij}$ be the number of times ward $i$ is judged to be superior to ward $j$ and $\boldsymbol{y}$ the set of outcomes for all pairs of wards. Let $M$ denote the number of pairs wards which were compared i.e. $M=\sum_{i=1}^{N}\sum_{j<i}\mathbf{1}(n_{ij}>0)$, where $\mathbf{1}(\cdot)$ denotes the indicator function. Under the reasonable assumption that comparisons between different pairs of wards and also comparisons for a given pair are independent, the likelihood of the observed data is a product of $M$ Bernoulli contributions
$\pi(\boldsymbol{y}|\boldsymbol{\lambda})=\prod_{k=1}^{M}\pi_{{ij}}^{y_{ij}}(1-\pi_{ij})^{n_{ij}-y_{ij}}\,,$ (2)
where in comparison $k$ ward $i$ was judged to be superior than ward $j$, and $\boldsymbol{\lambda}=\{\lambda_{1},\ldots,\lambda_{N}\}$ is the set of ward quality parameters.
@@ 2.2 The Bradley–Terry Model with Ties
We can relax the assumption that comparisons are Bernoulli trials and allow for tied comparisons. Examples of tied comparisons include tied matches in sporting events or when judges cannot distinguish between the qualities of two wards. We follow Rao and Kupper, (1967) and assume the probability ward is $i$ judged to have a higher rate of abuse than ward $j$ and a tie between wards $i$ and $j$ are
$\pi_{ij}$ $=\frac{\exp(\lambda_{i})}{\exp(\lambda_{i})+\exp(\lambda_{j}+\delta)},$ (3)
$\pi_{i=j}$ $=\frac{(\exp(2\delta)-1)\exp(\lambda_{i}+\lambda_{j})}{(\exp(\lambda_{i})+\exp(\lambda_{j}+\delta))(\exp(\lambda_{i}+\delta)+\exp(\lambda_{j}))}\quad(i\neq j,1\leq i,j\leq N).$ (4)
The parameter $\delta\geq 0$ contributes to the probability of a comparison resulting in a tie.
Unlike the standard model, each comparison in this model has one of three outcomes. As such, we must now consider all permutations, not combinations, of wards. We denote the the number of times ward $i$ was judged to have a higher rate of abuse than ward $j$ by $y_{ij}$ and the number of ties by $t_{ij}$, and the corresponding vectors by $\boldsymbol{y}$ and $\boldsymbol{t}$. We let $\boldsymbol{x}_{ij}$ be an $N\times 1$ vector with all elements being zero apart from the $i^{th}$ and $j^{th}$ elements which take values $1$ and $-1$ respectively. The likelihood function for this model is
$\pi(\boldsymbol{y},\boldsymbol{t}\mid\boldsymbol{\lambda},\delta)$ $\propto\prod_{i=1}^{N}\prod_{j=1}^{N}\pi_{ij}^{y_{ij}}\pi_{ji}^{y_{ji}}\pi_{i=j}^{t_{ij}/2}$
$\propto(\exp(2\delta)-1)^{\sum_{i=1}^{N}\sum_{j=1}^{N}{t_{ij}/2}}\prod_{i=1}^{N}\prod_{j=1}^{N}\pi_{ij}^{y_{ij}+t_{ij}}$
$\propto(\exp(2\delta)-1)^{\sum_{i=1}^{N}\sum_{j=1}^{N}{t_{ij}/2}}\prod_{i=1}^{N}\prod_{j=1}^{N}\left(\frac{\exp(\boldsymbol{x}_{ij}^{T}\boldsymbol{\lambda}-\delta)}{(1+\exp(\boldsymbol{x}_{ij}^{T}\boldsymbol{\lambda}-\delta))}\right)^{y_{ij}+t_{ij}}.$
@@ 3 Scalable Bayesian Inference for the Tied Model
Existing methods for drawing samples for this posterior distribution have several limitations. The method described in Caron and Doucet, (2012) uses an exponential latent variable representation of the model, which is an effective method for providing scalable inference. However, this method requires independent prior distributions to be placed on the ward parameters $\boldsymbol{\lambda}$, which is not suitable for our motivating example, where spatial correlation can be used to reduce the number of comparisons collected. The authors also use a flat, improper prior on the interval $[1,\infty]$ for $\exp(\delta)$. This results in a full conditional distribution that can be sampled from directly, but limits the inclusion of prior information about $\delta$. The Metropolis-Hasting method described in Seymour et al., 2022b could be extended to allow for tied comparisons and it allows for correlated prior distributions, reducing the number of comparisons that need to be collected. It would also have no restriction on the prior distribution for $\delta$. However, it scales poorly with the number of ward parameters $\boldsymbol{\lambda}$, and is slow to mix. In the example described in the paper, running the MCMC algorithm for 1,500,000 iterations requires 500,000 iterations to be removed as a burn-in period.
3.1 The Pólya-Gamma Latent Variable Representation
In the spirit of Caron and Doucet, (2012), we develop a latent variable representation of the BT model, but ensure that the representation allows for correlated prior distributions and does not restrict the choice of prior distribution on $\delta$. Consider the likelihood contribution from all comparisons, tied or otherwise, involving wards $i$ and $j$,
@@ 3.1 The Pólya-Gamma Latent Variable Representation
We introduce a latent variable $z_{ij}$ which follows a Pólya-Gamma (PG) distribution with density equal in distribution to the sum of gamma random variables $g_{k}\sim\hbox{Gamma}(b,1)$ given by
[…]
where $c\in\mathbb{R}$. Following Polson et al., (2013), by introducing a PG latent variable with parameters $b=1$ and $c=0$, we can write the likelihood contributions from the pair of wards $i$ and $j$ by
@@
Considering all pairs of wards observed in the study, the observed data likelihood function can be written as
$\pi(\boldsymbol{y},\boldsymbol{t}|\boldsymbol{\lambda},\delta)\propto$ $\int_{\boldsymbol{z}}\pi(\boldsymbol{y},\boldsymbol{t},\boldsymbol{z}|\boldsymbol{\lambda},\delta)\mbox{ d}\boldsymbol{z}$
$=$ $\prod_{i=1}^{N}\prod_{j=1}^{N}\pi(y_{ij},\,t_{ij}\mid\boldsymbol{\lambda},\delta)$
$\propto$ $\prod_{i=1}^{N}\prod_{j=1}^{N}\Biggl{(}\exp\left(\frac{1}{2}(\boldsymbol{x}_{ij}^{T}\,\boldsymbol{\lambda}-\delta)\right)$
@@
Including the latent variables $z_{ij}$ allow us to derive a closed form for the full conditional distribution for the ward parameters $\boldsymbol{\lambda}$, yielding a fast MCMC algorithm. We follow Seymour et al., 2022b a place a multivariate normal distribution on the ward parameters $\boldsymbol{\lambda}\sim N(\boldsymbol{\mu},\Sigma)$. This allows for prior correlation to be assumed between the ward parameters and the choice of $\Sigma$ is a modelling choice to suit the application.
@@
where the last step follows by completing the square. We can treat the double product as matrix multiplication, and denoting by $X$ the design matrix formed by $\boldsymbol{x}_{ij}^{T}$ for all pairs of $i$ and $j$, and by $\boldsymbol{\zeta}$ the vector containing $\delta-k_{ij}/z)_{ij}$, again for all pairs $i$ and $j$. The full conditional resolves to
@@
$\boldsymbol{\lambda}\mid\boldsymbol{y},\boldsymbol{t},\boldsymbol{z},\theta\sim N(\boldsymbol{\mu},S)$ (5)
where $S=\left(X^{T}ZX+\Sigma^{-1}\right)^{-1}$, $\boldsymbol{\mu}=SX^{T}\left((\boldsymbol{y}+\boldsymbol{t})/2-\theta Z\boldsymbol{1})\right)$, and $Z=\hbox{diag}(z_{ij})$.
The posterior distribution for $z_{ij}$ is
$z_{ij}\mid\boldsymbol{\lambda},\boldsymbol{y},\boldsymbol{t},\delta\sim PG(y_{ij}+t_{ij},\boldsymbol{x}_{ij}^{T}\boldsymbol{\lambda}-\delta).$
The choice of prior distribution on the tie parameter $\delta$ is a modelling choice and our method allows for any non-negative prior distribution to be placed on $\delta$. For our motivation, we place an exponential prior distribution with rate $\chi$ on $\delta$. The full conditional distribution has no closed form and so we generate samples for $\delta$ using a MHRW algorithm. The full conditional distribution is given by

### v2

@@ Abstract
Honour based abuse covers a wide range of family abuse including female genital mutilation and forced marriage. Safeguarding professionals need to identify where abuses are happening in their local community to best support those at risk of these crimes and take preventative action. However, there is little local data about these kinds of crime. To tackle this problem, we ran comparative judgement surveys to map abuses at local level, where participants where shown pairs of wards and asked which had a higher rate of honour based abuse. In previous comparative judgement studies, participants reported fatigue associated with comparisons between areas with similar levels of abuse. Allowing for tied comparisons reduces fatigue, but increase the computational complexity when fitting the model. We designed an efficient Markov Chain Monte Carlo algorithm to fit a model with ties, allowing for a wide range of prior distributions on the model parameters. Working with South Yorkshire Police and Oxford Against Cutting, we mapped the risk of honour based abuse at community level in two counties in the UK.
keywords:
Bradley–Terry Model, Bayesian Computation, Preference Learning, Violence Against Women and Girls
@@ 1 Introduction
Comparative judgement models have been to used to analyse a wide range of problems from predicting the outcomes of football leagues [4] to student assessment [7]. The aim of a comparative judgement study is to estimate the qualities of a set of objects. In a comparative judgement study, participants are shown pairs of objects and asked which meets object more strongly aligns with the study question. By allowing study participants to compare objects in terms of the quantity of interest rather than give absolute assessments of that quantity, comparative judgement models can elicit reliable and informative responses. In [11], the authors give the example of participants finding it easier to decide which of two circles has the larger area than estimating the size of an individual circle. The Bradley–Terry (BT) model [2] is perhaps the most common model for analysing comparative judgement data. It used to rank and estimate features of a set of objects based on pair-wise comparisons. It assigns a quality parameter to each object in the study and estimates the value of these parameters from pairwise comparisons.
@@
We were motivated by a project to tackle honour based abuse in the UK and support the development of community level safeguarding strategies. The specific aim of our project was to map the prevalence of honour based abuse at local level in two counties in the UK. Abuse to control family members in order to protect the family’s honour or standing is called honour based abuse. Perpetrators of honour based abuse often perceive a victim to be putting the family’s honour at risk, or to be bringing shame upon the family [1]. Examples include Female Genital Mutilation (FGM), a practice that involve partial or total removal of the external female genitalia for non-medical reasons (typically to evidence that the girl is a virgin on the wedding day, and as such is an honourable bride), or forced marriage, where individuals are coerced to marry someone who is perceived to be honourable. These abuse causes significant physical and emotional harm to the victim and can cause both immediate and long-term health issues.
@@
To map the prevalence of abuse at community level, we considered analysing existing data. However, we found existing data to be insufficient. Currently, the UK Government only provides regional level data about forced marriage, and national level data about other types of honour based abuse [6]. This makes developing effective safeguarding strategies for specific communities difficult. Other local level data does exist, as those providing services to victims may collect their own data. However, victims may be supported by a number of independent services, for example, law enforcement, health professionals or women’s refuges [1, §$\S 7$]. This means that no one service has the complete picture of where honour based abuse is happening. Accessing data from independent services would require a large number of data sharing agreements and the transfer of sensitive data about victims of abuse, which was not feasible. Thus, we decided to collect our data using a method that avoids collecting personally identifying information.
@@
We decided to carry out a comparative judgement study. A comparative judgement study works by showing study participants pairs of wards in the community and asking them to choose which ward has the higher rate of the abuse in question. Study participants, known as judges, should be people who are familiar with the community and abuse, however the wisdom of the crowd type approach means that judges do not need to have a perfect knowledge. Running a comparative judgement study is a viable and effective way to measure the rate of abuses at community level, without the need to collect personally identifying or sensitive data. Indeed, it has been previously used to map forced marriage at community level in Nottinghamshire [14] and deprivation at sub-ward level in Dar es Salaam, Tanzania [16]. However, in previous studies, judges reported fatigue and difficulty providing pairwise comparisons when the levels of abuse in pairs of wards were similar [14]. In particular judges, found it difficult to repeatedly make comparisons about pairs of wards which both had low levels of abuse. This has also been found to be an issue when using adaptive comparative judgement study designs [11]. To mitigate this issue in our study, we allowed the judges to say (i) the level of abuse in one ward was higher than the other, (ii) the level of abuse was similar – or tied – in both wards, or (iii) they wanted to skip that particular pair of wards.
The statistical aim of our work was to develop a method to fit a variant of the BT model that allows tied comparisons. In particular, we wanted to develop a method to fit such a model in a fast way, because our project partners had limited computing resources. Additionally, we wanted our method to allow for prior spatial assumptions to be included in the model. This is because in both counties there were a limited number of judges with the expertise to take part in our study, and including prior spatial assumptions has been shown to produce accurate results with a limited number of judges [16, 14].
Variants of the BT model allowing for tied comparisons are well developed, with the model described in [13] commonly used. The outcome of each comparison is modelled by a multinomial distribution, with the events being a win, a loss, and a tie. Inference for models with ties is, however, limited. Inference methods for simple models allowing for tied comparisons are fast and efficient but require large amounts of data. Inference methods have been developed for more sophisticated models can reduce the amount of data that need to be collected, but are slow to and expensive to use.
@@
We developed a new representation of the tied model by introducing a series of Pólya-Gamma latent variables. This allows for both sophisticated prior correlation between the model parameters and scalable inference. or computing maximum likelihood estimates, an iterative algorithm is developed in [13]. Maximum likelihood methods exist [8], but do not allow for prior spatial correlation to be assumed. Two Markov Chain Monte Carlo methods exist for generating samples from posterior distributions relevant to this model. In [3], the authors introduce a series of exponentially distribution latent variables to perform scalable inference. However, this method requires independent prior distributions to be placed on the model parameters, and an improper prior distribution on the parameter determining the importance of ties. In contrast, in [16], the authors propose a Metropolis-Hastings random walk algorithm that allows for a wide range of correlations between the model parameters, and a choice of any prior distribution on the tie parameter, but scales poorly with respect to the number of wards in the study.
@@ 1.1 Data
In the county of Oxfordshire, we worked with Oxford Against Cutting, a charity tackling FGM and honour based abuse in the county. We focused on the city of Oxford and town of Banbury in this study, and judges were shown pairs of wards in these two locations and asked which had a higher rate of honour based abuse. Judges could also say the wards had a similar level of abuse. We collected 766 comparisons, of which 199 (26.0%) were ties.
@@
1.2 Structure of the article
The remainder of the article is structured as follows. In Section 2, we describe the standard BT model and the variant of the model that allows for tied comparisons. We then describe an MCMC algorithm for fitting the the model with tied comparisons to data in Section 3. In Section 4, we carry out simulation studies to determine the sensitivity of our model to the choice of hyperparameters and the scalability of the MCMC algorithm. In Section 5, we use our MCMC algorithm to fit the model to analyse data we collected about honour based abuse in the counties of South Yorkshire and Oxfordshire, UK. In Section 6, we discuss our findings and method, and make suggestions about further work.
@@ 2.1 The Standard Bradley–Terry model
One of the most widely used models for modelling comparative judgement data is the-so-called BT model [2] and is defined as follows. Consider a set of $N$ wards. We assign a parameter $\lambda_{i}\in\mathbb{R}$ to each ward describing its quality. Large positive values of $\lambda_{i}$ correspond to wards with low prevalence of honour based abuse and large negative values to wards with high prevalence of abuse. When comparing wards $i$ and $j$, the probability ward $i$ is judged to be superior to ward $j$ (for $i\neq j$) depends on the difference in quality parameters
$\textrm{logit}(\pi(i\textrm{ beats }j))=\lambda_{i}-\lambda_{j}\iff\pi(i\textrm{ beats }j)=\frac{\exp(\lambda_{i})}{\exp(\lambda_{i})+\exp(\lambda_{j})}$ (1)
We can also write this probability in terms of the vector of regressors. Let $\boldsymbol{x}_{ij}$ be a vector of zeroes except the $i^{th}$ and $j^{th}$ entries, which take the values 1 and -1 respectively. By denoting the set of all ward quality parameters $\boldsymbol{\lambda}$, we have $\boldsymbol{x}^{T}_{ij}\boldsymbol{\lambda}=\lambda_{i}-\lambda_{j}$ and the probability can be written as
$\pi(i\textrm{ beats }j)=\frac{\exp\left(\boldsymbol{x}^{T}_{ij}\boldsymbol{\lambda}\right)}{1+\exp\left(\boldsymbol{x}^{T}_{ij}\boldsymbol{\lambda}\right)}.$ (2)
We will use the formulation in Equation (2) in the construction of our MCMC algorithm.
Under the assumption that comparisons between different pairs of wards and also comparisons for a given pair are independent, the likelihood of the observed data is the following product of Binomials trials for each pair of wards. Let $n_{ij}$ be the number of times wards $i$ and $j$ are compared, $y_{ij}$ be the number of times ward $i$ is judged to be superior to ward $j$ and $\boldsymbol{y}$ the set of outcomes for all pairs of wards. The likelihood function can be written as
$\pi(\boldsymbol{y}|\boldsymbol{\lambda})=\prod_{i=1}^{N}\prod_{j<i}\begin{pmatrix}n_{ij}\\ y_{ij}\end{pmatrix}\pi(i\textrm{ beats }j)^{y_{ij}}(1-\pi(i\textrm{ beats }j))^{n_{ij}-y_{ij}}\,,$ (3)
where $\boldsymbol{\lambda}=\{\lambda_{1},\ldots,\lambda_{N}\}$ is the set of ward quality parameters.
@@ 2.2 The Bradley–Terry model with ties
We can relax the assumption that comparisons are Bernoulli trials and allow for tied comparisons. Examples of tied comparisons include tied matches in sporting events or, in our motivating example, when judges cannot distinguish between the levels of honour based abuse in two wards. We follow [13] and assume the probability ward is $i$ judged to have a higher rate of abuse than ward $j$ and a tie between wards $i$ and $j$ (for $i\neq j$) are
$\pi(i\textrm{ beats }j)$ $=\frac{\exp(\lambda_{i})}{\exp(\lambda_{i})+\exp(\lambda_{j}+\delta)}$ (4)
$=\frac{\exp\left(\boldsymbol{x}^{T}_{ij}\boldsymbol{\lambda}-\delta\right)}{1+\exp\left(\boldsymbol{x}^{T}_{ij}\boldsymbol{\lambda}-\delta\right)},$ (5)
$\pi(i\textrm{ ties }j)$ $=\frac{(\exp(2\delta)-1)\exp(\lambda_{i}+\lambda_{j})}{(\exp(\lambda_{i})+\exp(\lambda_{j}+\delta))(\exp(\lambda_{i}+\delta)+\exp(\lambda_{j}))}.$ (6)
The parameter $\delta\geq 0$ contributes to the probability of a comparison resulting in a tie, with larger values of $\delta$ corresponding to a higher probability of a tie.
Unlike the standard model, each comparison in this model has one of three outcomes. As such, we must now consider all permutations, not combinations, of wards. We denote the the number of times ward $i$ was judged to have a higher rate of abuse than ward $j$ by $y_{ij}$ and the number of ties by $t_{ij}$, and the corresponding vectors by $\boldsymbol{y}$ and $\boldsymbol{t}$. The likelihood function for this model is
$\pi(\boldsymbol{y},\boldsymbol{t}\mid\boldsymbol{\lambda},\delta)$ $\propto\prod_{i=1}^{N}\prod_{j=1}^{N}\pi(i\textrm{ beats }j)^{y_{ij}}(1-\pi(i\textrm{ beats }j))^{y_{ji}}\pi(i\textrm{ ties }j)^{t_{ij}/2}$ (7)
$\propto(\exp(2\delta)-1)^{\sum_{i=1}^{N}\sum_{j=1}^{N}{t_{ij}/2}}\prod_{i=1}^{N}\prod_{j=1}^{N}\pi_{ij}^{y_{ij}+t_{ij}}$ (8)
$\propto(\exp(2\delta)-1)^{\sum_{i=1}^{N}\sum_{j=1}^{N}{t_{ij}/2}}\prod_{i=1}^{N}\prod_{j=1}^{N}\left(\frac{\exp(\boldsymbol{x}_{ij}^{T}\boldsymbol{\lambda}-\delta)}{(1+\exp(\boldsymbol{x}_{ij}^{T}\boldsymbol{\lambda}-\delta))}\right)^{y_{ij}+t_{ij}}.$ (9)
@@ 3 Scalable Bayesian inference for the tied model
An MCMC algorithm is needed to generate samples from this posterior distribution. As we are motivated by mapping honour-based abuse with local police forces and charities, it was important for us to develop an MCMC algorithm for the BT model with ties that was both allowed prior spatial correlation and was scalable. Prior spatial correlation in comparative judgement studies has been shown to considerably reduce the amount data that needs to be collected to obtain accurate estimates [16], which us why we wished to include it. Scalability was important due to the computing resources of our partners. Computing resources for fitting models in local police forces and charities are extremely limited and running an MCMC algorithm for multiple hours is likely infeasible. Our partners in this project used laptops with small memory allocations, which needed to be used for other day-to-day tasks like online meetings and word processing.
Currently, an MCMC algorithm does exist generate samples from the BT model with ties [3]. This algorithm includes an exponentially distributed latent random variable, which allows the MCMC algorithm to scale well as the number of wards increases. However, this algorithm requires the ward quality parameters to be a priori independent. Given the strong spatial component to our application, we wished to assume spatial prior correlation on these parameters. Additionally, the algorithm requires an improper prior distribution on the interval $(0,\infty)$ for the tie parameter $\delta$. This results in a full conditional distribution for $\delta$ that can be sampled from directly, but limits the inclusion of prior information about $\delta$
A Metropolis-Hastings random walk for the BT model without tied comparisons was developed in [16]. Although this does not allow for tied comparisons, it does allow to spatial prior correlation to be assumed for the ward quality parameters. This method could be extended to include inference for a tie parameter $\delta$ in a straightforward manner. However, the algorithm scales very poorly with respect to the number of wards. In the example used in [16], the authors use a Metropolis-Hastings random walk algorithm on a comparative judgement data set on deprivation in the 452 subwards of Dar es Salaam, Tanzania. Producing estimates for the quality parameters for all 452 subwards requires running the MCMC algorithm for 1,500,000 iterations. The Markov chains are slow to converge and their mixing is poor. Indeed, the first 500,000 iterations of the Markov chains are removed as a burn-in period, equating to around one hour of computer time. Extending this method to allow inference for a tie parameter would likely increase the computational cost of this method.
As the inference for the BT model with ties has shown to be scalable when a latent variable is included [3], we decided to develop a latent variable representation of the BT model with ties, but ensure that the representation allows for correlated prior distributions and does not restrict the choice of prior distribution on $\delta$. For each pair of wards, we introduce a latent variable $z_{ij}$, which follows a Pólya-Gamma (PG) distribution. This distribution has density equal in distribution to the sum of gamma random variables $g_{k}\sim\hbox{Gamma}(b,1)$ given by
[…]
where $c\in\mathbb{R}$. In [12], the authors show the PG distribution has the useful property
[…]
As contributions to our likelihood function have the same functional form as the left hand side of Equation (10), we will use this property to derive the latent variable augmented likelihood function. Consider the likelihood contribution from all comparisons, tied or otherwise, involving wards $i$ and $j$ (that is the $(i,j)^{th}$ term of the observed data likelihood function in Equation 9),
@@
Using the property in Equation (10), we can write this likelihood contribution as
@@
The likelihood function can then be written as the product of this likelihood contribution over all pairs of wards, yielding
$\pi(\boldsymbol{y},\boldsymbol{t}|\boldsymbol{\lambda},\boldsymbol{z},\delta)\propto$ $\prod_{i=1}^{N}\prod_{j=1}^{N}\pi(y_{ij},\,t_{ij}\mid\boldsymbol{\lambda},z_{ij},\delta)$
$\propto$ $\,(\exp(2\delta)-1)^{\sum_{i=1}^{N}\sum_{j=1}^{N}{t_{ij}/2}}\prod_{i=1}^{N}\prod_{j=1}^{N}\Biggl{(}\exp\left(\frac{1}{2}(\boldsymbol{x}_{ij}^{T}\,\boldsymbol{\lambda}-\delta)\right)$
@@
By Bayes’ theorem, this posterior distribution including the latent variables $\boldsymbol{z}$ is
$\pi(\boldsymbol{\lambda},\boldsymbol{z},\delta\mid\boldsymbol{y},\boldsymbol{t})\propto\pi(\boldsymbol{y},\boldsymbol{t}|\boldsymbol{\lambda},\boldsymbol{z},\delta)\pi(\boldsymbol{\lambda})\pi(\boldsymbol{z})\pi(\delta).$
We will now show that when a joint prior distribution is placed on the ward parameter $\boldsymbol{\lambda}$, the inclusion of the latent variables $z_{ij}$ allow us to derive a closed form for the full conditional distribution for $\boldsymbol{\lambda}$. The result is a scalable MCMC algorithm for the BT model with ties including prior spatial correlation.
To include prior correlation in the BT model, we follow [16] and place a multivariate normal distribution on the ward parameters $\boldsymbol{\lambda}\sim N(\boldsymbol{\mu},\Sigma)$. This allows for prior correlation to be assumed between the ward parameters and the choice of $\Sigma$ is a modelling choice to suit the application. In our application, we choose a network based spatial structure, where each of the wards are nodes and edges are placed between adjacent wards. The full conditional distribution for the ward parameters using this prior distribution is
@@
where $k_{ij}=y_{ij}-n_{ij}/2$. The last step follows by completing the square. We can treat the double product as matrix multiplication, and denoting by $X$ the design matrix where each column is the vector of regressors $\boldsymbol{x}_{ij}^{T}$, and by $\boldsymbol{\zeta}$ the vector containing $\delta-k_{ij}/z_{ij}$, for all pairs $i$ and $j$. The full conditional distribution resolves to
@@
$\boldsymbol{\lambda}\mid\boldsymbol{y},\boldsymbol{t},\boldsymbol{z},\delta\sim N(\boldsymbol{\mu},S)$ (11)
where $S=\left(X^{T}ZX+\Sigma^{-1}\right)^{-1}$, $\boldsymbol{\mu}=SX^{T}\left((\boldsymbol{y}+\boldsymbol{t})/2-\delta Z\boldsymbol{1})\right)$, and $Z=\hbox{diag}(z_{ij})$.
The full conditional distribution for $z_{ij}$ is
$z_{ij}\mid\boldsymbol{\lambda},\boldsymbol{y},\boldsymbol{t},\delta\sim PG\left(y_{ij}+t_{ij},\,\boldsymbol{x}_{ij}^{T}\boldsymbol{\lambda}-\delta\right).$
Sampling from this distribution can be efficiently done using an accept-reject algorithm based on the alternating-series method of [5], and we use the function provided in the BayesLogit R package [12].
The choice of prior distribution on the tie parameter $\delta$ is a modelling choice and our method allows for any non-negative prior distribution to be placed on $\delta$. For our motivation, we place an vague exponential prior distribution with rate $\chi$ on $\delta$. The full conditional distribution has no closed form and so we generate samples for $\delta$ using a Metropolis-Hastings random walk algorithm. The full conditional distribution is given by